OpenStudy (baby456):

6.Jane is baking cookies for a large party.She has a recipe that will make one batch of two dozen cookies,and she decides to make five batches. To make five batches,she finds that she will need 12.5 cups of flour and 15 eggs. Write an equation to describe each of the following situations. then solve the probelm. a. how many cookies will she make in all?

OpenStudy (texaschic101):

1b (or just b) = 24 cookies (2 dozen) 5b = 5(24) = 120 cookies (10 dozen) 1 batch makes 24 cookies, so 5 batches makes 120 cookies.(or 10 dozen) For 5 batches, she needs 12.5 cups of four and 15 eggs For 1 batch, she needs : 12.5/5 = 2.5 (or 2 1/2) cups of flour 15/5 = 3 eggs I am sorry that I didn't really make any equations.... @ganeshie8 ..what would the equations be ?

ganeshie8 (ganeshie8):

for a, i wudnt make any equations either...

ganeshie8 (ganeshie8):

if they insist for equations, d = 12 b = 2d c = 5b = 5(2d) = 5(2(12)) =120

OpenStudy (texaschic101):

thats works